• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name mercado

    Log entry time 08:03:30 on October 2,2009

    Entry number 292716

    keyword=Owl Shift Summary

    Shift Summary: Owl Shift, Oct 2, '09

    Shift started with DAQ problems. DAQ was rejecting over half
    the data because of helicity sequence errors. Called Bob Michaels
    in for assistance.

    While he got here, I went to the Injector and rebooted the crates.

    I also looked at some runs to try and diagnose the problem. From the data,
    it seemed that signals for helicity and pairsynch looked the same. Looked
    at the signals on a scope and they indeed looked similar.

    Went to MCC with Bob to check on the helicity control board settings.

    Called Kent to verify the settings were correct.

    Changed back from 'Output Select'=Helicity Delay to 'Output Select'=Pair Sync.

    Started run 12876 at 1:21 and the data looks good now.

    Taking production data...

    2:19 - Run 12877

    5:03 - MCC called - They reported a 'major' problem. They are calling in an expert.

    5:23 - MCC called - They are waiting for experts to get here in order to fix an RF problem in the
    North Linac. They say it will be a while until beam comes back.

    6:15 - MCC called - They will try to deliver beam again.

    6:18 - Once again at 100uA.

    6:22 - MCC called - They are having the same problem again.

    8:00 - Still no beam...